OpenStudy (anonymous):

A pencil at a stationery store costs $1, and a pen costs $1.50. Shawn spent $12 at the store. He bought a total of 9 items. Which system of equations can be used to find the number of pencils (x) and pens (y) he bought? 1.5x + y = 12 x = 9y x + 1.5y = 12 x + y = 9 x + 9y = 12 x = 1.5y 9x + y = 12 x = 1.5y

OpenStudy (anonymous):

x + 1.5y = 12 x + y = 9

OpenStudy (baby_bear69):

x stands for how many pencils there are for $1 and y stands for how many pens there are so $1.5. Together they costed $12 so x+1.5y=12 There were 9 pencils and pens together, so x+y = 9

OpenStudy (anonymous):

Look for the option that has the TOTAL number of pencils and pens. Pencils mean x and pens mean y. He bought a total of 9. So, x + y = 9. Finding this will rule out all other options